OpenStudy (anonymous):

A spinner numbered 1 through 10 is spun 3 times. What is the probability of spinning a 5, then an even number, and then another even number? Write your answer as a decimal. Do not round.

OpenStudy (anonymous):

So I know spinning a 5 is 1/10 Then spinning an even number is 5/10=1/2 Then another even number is 5/10=1/2 How would I solve this?

OpenStudy (anonymous):

@iambatman Help please?